# FinTSBridge: A New Evaluation Suite for Real-world Financial Prediction with Advanced Time Series Models

### v1

@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1.1 Discription
The GSMI dataset includes 20 major indices from global stock markets, recording daily price and trading volume data for these indices over nearly 20 years, from 2005 to 2024. We have introduced each index in the table below. Index prediction has always been an important and challenging issue, whether it is forecasting the future price trends of the indices or predicting market trading volumes, both of which have significant implications for investment and trading behavior. Research has shown that there are time-varying and complex relationships between global market indices (Youssef et al., 2021; Baltussen et al., 2019; Boyer et al., 2006; Wu et al., 2021b). These complex interrelationships between indices highlight the challenges of forecasting using global index datasets. By leveraging state-of-the-art time-series AI models, we aim to explore the frontier of this important issue and provide a more challenging dataset for time-series prediction, which will better demonstrate the performance differences between different models.
@@ B.1 Configurations
Due to the large number of models and tasks involved, we were unable to extensively tune hyperparameters to find the best configuration for every model. However, hyperparameter tuning was not the primary focus of this work. In these experiments, PSformer is from Wang et al., and the other models are from TSLib (Wang et al., 2024b). Regarding hyperparameters, if a model has corresponding Exchange script configurations, we will adopt those settings as the Exchange dataset is more aligned with the characteristics of our financial datasets. Otherwise, we will use the script parameters from ETT. All the experiments are conducted on NVIDIA A100 80GB GPU.
@@ B.5 Naive Model Details
Forecast Samples. The Figure 11 illustrates the prediction samples of this method across different datasets. The yellow line represents the Naive prediction, which remains nearly constant and fails to capture the dynamic variations in the time series. However, for time series without significant periodicity, the Naive model may achieve relatively small prediction errors. Performance Comparison. The Table 2 presents the MSE and MAE losses of the Naive model on the Exchange, ETTh2, ETTm2, and Traffic datasets. For comparative analysis, the results of three other time series models (iTransformer, DLinear, and FEDformer results from (Liu et al., 2024)) are also included. The results suggest that the Naive model performs poorly on time series with strong periodicity, such as Traffic, but achieves smaller prediction errors on datasets like Exchange. This further emphasizes that prediction error alone is insufficient to comprehensively evaluate the forecasting capability of time series models.

### v2

@@ Appendix A Dataset Details
Table 3: Summary of new datasets.
Dataset GSMI OPTION BTCF
Range (2000-01, 2024-12) (2024-04, 2024-12) (2020-01, 2024-11)
Variate 100 22 12
Samples 6533 37431 43014
Frequency Daily Minutely Hourly
Finance Aera Stock Indices Option Future & Spot
Predict Length {5,21,63,126} {5,21,63,126} {5,21,63,126}
@@ A.1.1 Discription
The GSMI dataset includes 20 major indices from global stock markets, recording daily price and trading volume data for these indices over nearly 20 years, from 2005 to 2024. We have introduced each index in the table below. Index prediction has always been an important and challenging issue, whether it is forecasting the future price trends of the indices or predicting market trading volumes, both of which have significant implications for investment and trading behavior. Research has shown that there are time-varying and complex relationships between global market indices Youssef et al. (2021); Baltussen et al. (2019); Boyer et al. (2006); Wu et al. (2021b). These complex interrelationships between indices highlight the challenges of forecasting using global index datasets. By leveraging state-of-the-art time-series AI models, we aim to explore the frontier of this important issue and provide a more challenging dataset for time-series prediction, which will better demonstrate the performance differences between different models.
@@ B.1 Configurations
Due to the large number of models and tasks involved, we were unable to extensively tune hyperparameters to find the best configuration for every model. However, hyperparameter tuning was not the primary focus of this work. In these experiments, PSformer is from Wang et al., and the other models are from TSLib Wang et al. (2024c). Regarding hyperparameters, if a model has corresponding Exchange script configurations, we will adopt those settings as the Exchange dataset is more aligned with the characteristics of our financial datasets. Otherwise, we will use the script parameters from ETT. All the experiments are conducted on NVIDIA A100 80GB GPU.
@@ B.5 Naive Model Details
Forecast Samples. The Figure 11 illustrates the prediction samples of this method across different datasets. The yellow line represents the Naive prediction, which remains nearly constant and fails to capture the dynamic variations in the time series. However, for time series without significant periodicity, the Naive model may achieve relatively small prediction errors.
Performance Comparison. The Table 13 presents the MSE and MAE losses of the Naive model on the Exchange, ETTh2, ETTm2, and Traffic datasets. For comparative analysis, the results of three other time series models (iTransformer, DLinear, and FEDformer results from Liu et al. (2024)) are also included. The results suggest that the Naive model performs poorly on time series with strong periodicity, such as Traffic, but achieves smaller prediction errors on datasets like Exchange. This further emphasizes that prediction error alone is insufficient to comprehensively evaluate the forecasting capability of time series models.
@@
Dataset Length Naive iTransformer DLinear FEDformer
MSE MAE MSE MAE MSE MAE MSE MAE
Exchange 96 0.081 0.196 0.086 0.206 0.088 0.218 0.148 0.278
192 0.167 0.289 0.177 0.299 0.176 0.315 0.271 0.315
336 0.306 0.398 0.331 0.417 0.313 0.427 0.460 0.427
720 0.810 0.676 0.847 0.691 0.839 0.695 1.195 0.695
ETTh2 96 0.281 0.369 0.297 0.349 0.333 0.387 0.358 0.397
192 0.339 0.404 0.380 0.400 0.477 0.476 0.429 0.439
336 0.372 0.424 0.428 0.432 0.594 0.541 0.496 0.487
720 0.439 0.463 0.427 0.445 0.831 0.657 0.463 0.474
ETTm2 96 0.202 0.310 0.180 0.264 0.193 0.292 0.203 0.287
192 0.235 0.337 0.250 0.309 0.284 0.362 0.269 0.328
336 0.270 0.361 0.311 0.348 0.369 0.427 0.325 0.366
720 0.335 0.401 0.412 0.407 0.554 0.522 0.421 0.415
Traffic 96 2.715 1.077 0.395 0.268 0.650 0.396 0.587 0.366
192 2.747 1.085 0.417 0.276 0.598 0.370 0.604 0.373
336 2.789 1.094 0.433 0.283 0.605 0.373 0.621 0.383
720 2.810 1.097 0.467 0.302 0.645 0.394 0.626 0.382